OpenStudy (anonymous):

A rectangular paperboard 35in long and 22in wide has a semicircle cut out of it, as shown below. Find the area of the paperboard that remains. Use the value 3.14 for π , and do not round your answer. Be sure to include the correct unit in your answer.

Directrix (directrix):

>> as shown below. Please post the diagram. Thanks.

OpenStudy (anonymous):

Ill screen shot it give me one second

Directrix (directrix):

The area of the rectangle minus 1/2 the area of a circle of radius 11 is what is left of the paperboard.

Directrix (directrix):

Area of rectangle is 35 * 22 = ? @heythatsness

OpenStudy (anonymous):

770

Directrix (directrix):

.5 * 3.14 * 11^2 = ? @heythatsness

OpenStudy (anonymous):

121

Directrix (directrix):

Copy and paste this into the Google search box and see if the Google calculator gets that: .5 * 3.14 * 11^2 =

Directrix (directrix):

Post what you get. @heythatsness

Directrix (directrix):

@heythatsness 121 is 11 times 11. Multiply that by one-half of 3.14

OpenStudy (anonymous):

189.97 is the correct answer fot that equation i did my math wrong

Directrix (directrix):

Subtract 189.97 (area of semicircle) from 770 which is the area of the rectangle to get the area of the remaining paper. 770 - 189.97 = ?

OpenStudy (anonymous):

solve it @heythatsness

Directrix (directrix):

>>> do not round your answer. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

580.03

Directrix (directrix):

square inches. Here is the calculation we did: 35 * 22 - ( .5 * 3.14 * 11^2) =

Directrix (directrix):

Question?